OpenStudy (anonymous):

Find the first derivative of f(θ)=(θ+1)cosθ

OpenStudy (amistre64):

replace that theta thing with a normal looking variable for starters :)

OpenStudy (amistre64):

(x+1)cos(x) distribute the cos thru x cos(x) + cos(x) now derive both of them seperately

OpenStudy (amistre64):

Dx(x cosx) + Dx(cosx)

OpenStudy (amistre64):

the first term is a product, so use the product rule on it...

OpenStudy (amistre64):

x Dx(cosx) + Dx(x) cosx + Dx(cosx)

OpenStudy (amistre64):

-x sin(x) + cos(x) - sin(x)

OpenStudy (nowhereman):

That looks rather complicated. I would do it like this: \[Df = \cos θ D(θ+1) + (θ+1)D\cos θ = \cos θ - (θ+1)\sin θ\]

OpenStudy (anonymous):

could both answers be used as the correct answer?

OpenStudy (nowhereman):

Yes, and they yield the same result after all.

OpenStudy (amistre64):

Both answers are exactly the same :) I just didnt factor mine out.

OpenStudy (anonymous):

oh well thank you so much :)